# 22affiliationtext: Booth School of Business

### v1

@@ D.2.1 Additional examples of concept composition
Figure D.2.1 contains additional examples of composing concepts from the NVAE-based context module. Each row corresponds to actual generated samples from the trained model in different contexts:
@@
For MNIST and 3DIdent, the final row of compositional generations is genuinely OOD in that the training data does not contain any examples where both concepts have been intervened upon. Recall that CelebA is trained using conditional as opposed to interventional samples as an ablation on the sensitivity to non-interventional data.
Additional examples of both single- and multi-concept interventions over more contexts are provided in Appendices D.2.2D.2.4.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 8: Additional examples of concept composition in MNIST (left), 3DIdent (middle), and CelebA (right). In MNIST and 3DIdent, the samples are OOD: The training data did not contain any examples with these concepts composed together. The CelebA results are an ablation to understand the effect of conditioning vs intervention, and so there is some leakage between concepts, where the interventional datasets (MNIST, CelebA) show no leakage.
D.2.2 Additional MNIST figures
Figure 9: Examples of single-node concept interventions in MNIST. (top) observational, scaled, shear (bottom) thic, thin, swel.
Figure 10: Examples of OOD multi-node concept interventions in MNIST.
D.2.3 Additional 3DIdent figures
[Uncaptioned image][Uncaptioned image]
Figure 11: Examples of single-node concept interventions in 3DIdent. (top) observational, background (bottom) object, spotlight.
Figure 12: Examples of OOD multi-node concept interventions in 3DIdent.
D.2.4 Additional CelebA figures
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 13: Examples of single-node concept interventions in CelebA. (top) observational, bags under eyes, black hair (bottom) blond hair, male, mouth slightly open.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 14: Examples of OOD multi-node concept interventions in CelebA.

### v2

@@ E.2.1 Additional examples of concept composition
Figure 15 contains additional examples of composing concepts from the NVAE-based context module. Each row corresponds to actual generated samples from the trained model in different contexts:
@@
The final row of compositional generations is genuinely OOD, in that the training data does not contain any examples where both concepts have been intervened upon.
Additional examples of both concept learning (sampling after single-target interventions) and composition (sampling after multi-target interventions) are provided in Appendices E.2.2E.2.3.
Figure 15: Additional examples of concept composition in MNIST (left) and 3DIdent (right).
E.2.2 Additional MNIST figures
See Figure 16 for concept learning and Figure 17 for composition.
Figure 16: Examples of learned concepts (sampling after single-target intervention) in MNIST. (top) observational, scaled, shear (bottom) thic, thin, swel.
Figure 17: Examples of OOD concept composition (sampling after multi-target intervention) in MNIST.
E.2.3 Additional 3DIdent figures
See Figure 18 for concept learning and Figure 19 for composition.
Figure 18: Examples of learned concepts (sampling after single-target intervention) in 3DIdent. (top) observational, background (bottom) object, spotlight.
Figure 19: Examples of OOD concept composition (sampling after multi-target intervention) in 3DIdent.